### v1

# Performance-lossless Black-box Model Watermarking

Abstract.
With the development of deep learning, high-value and high-cost models have become valuable assets, and related intellectual property protection technologies have become a hot topic. However, existing model watermarking work in black-box scenarios mainly originates from training-based backdoor methods, which probably degrade original task performance. To address this, we propose a branch backdoor-based model watermarking protocol to protect model intellectual property, where a construction based on a message authentication scheme is adopted as the branch indicator. We prove the lossless performance of the protocol by reduction. Taking the language generation task as an instance, we show the effectiveness of the proposed protocol.
performance-lossless; black-box; watermark; branch
1. Introduction
Recently, the development of deep learning technology represented by GPTs (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023b), GANs (Goodfellow et al., 2014; Brock et al., 2019; Karras et al., 2021), Diffusion models (Dhariwal and Nichol, 2021; Ho et al., 2022; Rombach et al., 2022), etc. has become increasingly mature. Applications (OpenAI, 2023a; Google, 2023; Anthropic, 2023) in Computer Vision (CV), Natural Language Processing (NLP), etc., have demonstrated tremendous commercial value. However, developing high-performance models entails substantial costs, including the involvement of machine learning experts with rich experience, extensive and high-quality task datasets, and expensive computational resources. Development teams maintain legitimate intellectual property rights over their produced models and often refrain from open-sourcing critical model details. They provide on-demand model functionality services to users through a commercial paradigm of remote access to model input-output APIs, such as OpenAI’s ChatGPT (OpenAI, 2023a), Google’s Bard (Google, 2023), and Anthropic’s Claude (Anthropic, 2023).
While development teams take great care in deploying these valuable model assets, they cannot completely prevent malicious users from disregarding intellectual property. We consider a threat scenario wherein model service providers (model development teams) possessing valuable model assets offer business arrangements allowing third-party application developers (application service providers) to access model functionality via black-box APIs. However, dishonest application developers may attempt to infringe on the intellectual property of model service providers. They do so by leveraging network access restrictions between different countries to resell API access rights to users, using low-quality models to impersonate high-performance models when deploying vertical applications such as chatbots to deceive uninformed application users, or even directly deploying the unlicensed model in the name of a self-developed high-performance model. Such behavior obtaining improper benefits has seriously damaged the intellectual property rights, reputation, and economic interests of the model development team. Security measures need to be taken to detect and confirm the legitimacy of the model.
Watermarking serves as an effective solution for protecting intellectual property. Existing model watermark works can generally be divided into three types based on the model knowledge needed for copyright verification or the ability to control the model: white-box watermarks, black-box watermarks, and no-box watermarks. White-box verification watermarking assumes that the model is completely transparent, allowing the verifier to manipulate the model within the verification rules to detect the watermark signal. In black-box verification watermarking scenarios, the verifier lacks knowledge of the inner workings of the model, but can manipulate the inputs of the model to obtain corresponding outputs to detect the watermark signal. In the case of no-box validation watermarks, the verifier only has access to the model’s output to detect the watermark signal. In the threat scenario described previously, suspect models are difficult to obtain, which limits white-box model watermarking methods. No-box model watermarking requires independently meaningful model outputs, which makes it useful only for generative tasks. Therefore, black-box model watermarking is more common and practical.
Black-box model watermarking works mainly rely on backdoor techniques. In 2018, Adi et al. (Adi et al., 2018) first established the connection between backdoor models and model watermarks, introducing the first backdoor-based model watermark work. Verifiers activate a pre-embedded backdoor in the model by inputting trigger samples during the inference process and detecting watermark signals in the output for model ownership confirmation. Existing backdoor-based model watermarks (Darvish Rouhani et al., 2019; Le Merrer et al., 2020; Adi et al., 2018; Chen et al., 2019; Li et al., 2022b; Jia et al., 2022; Lounici et al., 2021; Lao et al., 2022a; Jia et al., 2021; Guo and Potkonjak, 2018; Zhu et al., 2020; Maung Maung and Kiya, 2021; Li et al., 2022a, 2019; Lao et al., 2022b; Yang et al., 2021; Charette et al., 2022; Namba and Sakuma, 2019; Bansal et al., 2022; Zhong et al., 2020; Mehta et al., 2022) usually involve fine-tuning or retraining the model’s weights to embed the backdoor, which essentially constitutes multitask learning. Due to the incomprehensible nature of neural networks, despite attempts to separate the influences between the two tasks through carefully designed loss functions (Jia et al., 2021), weight selections (Lao et al., 2022a; Namba and Sakuma, 2019; Bansal et al., 2022), trigger sets (Darvish Rouhani et al., 2019; Adi et al., 2018; Zhang et al., 2018; Jia et al., 2022; Lounici et al., 2021; Lao et al., 2022a; Jia et al., 2021; Guo and Potkonjak, 2018; Zhu et al., 2020; Maung Maung and Kiya, 2021; Zhong et al., 2020), etc., there is no reliable theoretical analysis to prove that the watermarking task will not distort the normal task. Their experimental results also indicate that embedding watermarks impacts the accuracy of image classification models.
An ideal model watermark should not affect the model’s original functionality. In order to meet this requirement, we propose a performance-lossless black-box model watermarking, with specific contributions as follows:
  1. (1)
    We propose a branch watermarking protocol for black-box model copyright verification scenarios, which is a watermark embedding paradigm during the model service deployment phase. The protocol decouples the original task of the model from the verification task by inserting an independent watermark module in parallel to the target model within the API.
  2. (2)
    To achieve the watermark module’s independence, we propose the concept of the unlearnable trigger based on the provably unforgeable Message Authentication Code (MAC), and our constructions enjoy a formal security analysis that reduces the performance-lossless property of the model watermarking to the properties of the underlying cryptographic primitive.
  3. (3)
    Taking the language generation model as an example, we provide the instantiation strategy of the proposed watermark protocol, and combine it with the existing no-box watermarking scheme to improve the security against interference attacks.
2. Related Work
2.1. Backdoor Attacks
Backdoor attacks aim to embed hidden backdoors into deep neural networks, allowing the model to perform normally on regular samples but activate the internal backdoor when exposed to predefined triggers, resulting in malicious or poor-quality predictions. Depending on whether backdoor implantation involves training a victim model, these attacks are mainly divided into two types: training-based backdoors and training-free backdoors.
2.1.1. Training-based Backdoors
We can think of the initial training phase of the model and the fine-tuning phase of subsequent downstream tasks as the development phase of the model. Most backdoor attacks that occur during the development phase fall into the training-based backdoor category, whether by poisoning the training dataset or injecting malicious code into the training code.
Backdoor attacks based on poisoned data (Gu et al., 2017; Turner et al., 2019; Chen et al., 2017; Saha et al., 2020; Turner et al., 2018) have long been a topic of concern. When model developers face resource constraints, whether due to a lack of training data or limited computing power, they may purchase datasets from third parties or delegate training tasks to third parties, thereby exposing themselves to the risk of attackers injecting poisoned data into the training data. Attackers poison the training data, causing the model to learn the knowledge of the poisoned data during the training process to plant the backdoor. Backdoor attacks based on poisoned training code (Bagdasaryan and Shmatikov, 2021) constitute a form of blind attack on unknown models. Attackers insert malicious code into the model’s training code, automatically poisons the data and controls the loss function, and injects new tasks into the model during the training process to plant backdoors.
The knowledge associated with these poisoned data often leads to poor predictions. However, this backdoor technique can also be used for model copyright verification when the knowledge is used as a signature of model developers.
2.1.2. Training-free Backdoors
Compared with training-based backdoor methods, training-free backdoor methods inject backdoors completely independently of the training process of the model. Attackers insert backdoors by injecting poisonous branches into the model’s inference process during the model deployment phase, such as poisoning model structures or control flow. This type of method does not require secondary modification of the model weights, thereby avoiding damaging the original performance of the model.
Backdoor attacks based on the poisoned model structure (Liu et al., 2022; Zou et al., 2018; Tang et al., 2020; Bober-Irizar et al., 2023) involve attackers adding additional subnetworks to the victim model’s network structure to inject a backdoor. It is worth noting that the poisoned structure is only used during the inference process and will not affect the training process. Backdoor attacks based on poisoned control flow (Li et al., 2021; Clifford et al., 2022; Wang, 2022) occur when the victim model is deployed and run as part of software. Attackers insert new node modules into the control flow to inject a backdoor. Both attack types essentially involve inserting a branch into the victim model’s inference process, with the difference lying in the form of the backdoor trigger detector, whether it is a neural network or a diverse structure.
The mentioned attacks, particularly control flow poisoning, resemble traditional software attack techniques such as Trojans. The principles for designing an ideal Trojan horse attack, as proposed in the work (Tang et al., 2020), provide guidance for designing our branch watermarking protocol: 1) The Trojan horse attack should be model-agnostic, meaning it can be applied to different DNNs with minimal effort; 2) Inserting a Trojan horse in the model should not change the performance on the original task; 3) The triggers should be diverse; 4) The Trojan horse should be invisible and not detectable by existing Trojan horse detection algorithms.
2.2. Watermarking for Models
Model watermarking is a technology used to protect and track machine learning models, similar to the application of digital watermarks in digital media. The main goal of model watermarking is to embed a unique identifier or label in the model without affecting the performance or accuracy of the model, thereby identifying and verifying the legitimacy of the model. This helps prevent unauthorized copying, misuse or theft of the model while also serving as a means of proving the provenance and integrity of the model. In the context of the remote serving paradigm of models, we mainly focus on two types of model watermarks: black-box watermarks and no-box watermarks.
2.2.1. Black-box Watermarking
The “black box” is an image description of the verifier’s knowledge of the model, indicating that the verifier’s knowledge of the model is limited to queries and responses. Model owners use techniques such as backdoor attacks to embed signature-related watermark information in the model. They use specific trigger query-response pairs to detect watermarks for verification.
The mentioned Adi et al.’s work (Adi et al., 2018) introduced a relationship between malicious backdoor attacks and benign watermark defenses for the model, reducing the watermark task of deep neural networks to a backdoor task. Existing model watermarking approaches based on backdoor design share the same thought as methods such as training-based backdoors. Model owners use specific input-output pairs as watermark patterns, which is a form of knowledge that needs to be learned. This knowledge is injected into the model and trained to create backdoors as watermarks. Then, during ownership disputes, specific input triggers are used to activate the watermark backdoor and obtain relevant specific outputs for copyright verification. However, as discussed in section 2.1.2, these backdoor attacks rely on injecting the backdoor during training. The backdoor task and the model’s original task are entangled during the training process, which makes it challenging to ensure that the original task performance is not distorted. The same goes for watermark tasks.
2.2.2. No-box Watermarking
There is an ownership dispute scenario where the validator does not even have access to the “box” of the suspect model. The model owner embeds a special watermark identifier in the model’s output. When ownership disputes arise, under the strict condition that there are only some unknown inputs to the model, they can extract copyright information from the output to complete verification of the source model.
For tasks related to copyright protection of language generation models, methods such as synonym replacement-based no-box watermarking were proposed, and subsequent work further improved the watermark scheme’s impact on text quality (He et al., 2022b, a). Kirchenbauer et al. (Kirchenbauer et al., 2023) proposed a zero-bit watermark scheme that adds biased signals based on blacklists and whitelists to the outputs (i.e., logits) of the model. While the paper targeted a binary classification traceability task for human-machine text, it can be applied to model copyright verification tasks in a no-box scenario. Some works (Lee et al., 2023; Zhao et al., 2023a; Fu et al., 2023; Liu et al., 2023; Yoo et al., 2023) have extended and improved upon this work (Kirchenbauer et al., 2023). Zhao et al. (Zhao et al., 2023b) transformed the model watermark work (Charette et al., 2022) designed for image classification tasks, to language generation tasks in the context of no-box watermarking.
3. A Performance-lossless Branch Watermarking Protocol for Model
In this section, we provide a comprehensive description of the threat model targeted in this paper. We will then propose the corresponding model watermarking protocol and analyze the potential attacks that the watermarking protocol may face. Finally, we summarize the threat models and attack methods and outline the principles that watermarking protocols should satisfy in practical applications.
Figure 1. There are three roles in the threat scenario. The model development team provide black-box deployment to the application service provider. The latter may dishonestly provide application services to end users, which we summarize as two cases of dishonesty in the red zone.
3.1. Threat Model
In the threat model of the model remote service scenario, there are three key roles: the model development team, application service provider and end users, as shown in Figure 1.
The model development team consists of machine learning experts. They invest considerable resources and effort into developing high-performance models, such as large language models. They act as model owners and provide remote access to model functionality via APIs to application service providers in exchange for a fee.
The application service provider is a team of non-machine learning experts who purchase model functionality services from the model owner. They leverage remote APIs to deploy specific applications (such as chatbots) and deliver well-encapsulated, user-friendly, and easy-to-use model applications to end-users.
End users are customers who purchase specific application functionalities from the application service provider. They can enjoy model services without having relevant expertise.
The threats considered here come from dishonest application service providers. Honest application service providers will strictly abide by the API usage agreement, while dishonest application service providers may abuse the model’s remote API for unauthorized application development. Two situations are given here:
  1. (1)
    The application provider claims to use a well-known high-performance model A via the remote API interface but in fact uses a low-cost model B. Their aim is to profit from users at low cost while damaging the reputation of the model in question.
  2. (2)
    The application provider uses a well-known high-performance model A through the remote API but deliberately claims that it is a self-developed model B. This allows them to gain market share and profits while compromising the intellectual property of model owners.
In practical situations, the model owner and the application service provider may be the same team. However, this involves some internal personnel management issues that are not within the scope of this research. This paper treats the model owner and application service provider as two independent entities and primarily explores external threats to the model owner.
3.2. Planting A Parallel Branch in Model API
Remote services have the advantages of a black box model, allowing watermark modules to be added to the API. We propose a branch backdoor-based watermarking protocol for model. The idea of the mutual transformation between malicious backdoors and benign watermarks initially originated from Adi et al.’s work (Adi et al., 2018), but unlike previous work, we use a training-free backdoor that modifies the control flow rather than a training-based backdoor. The advantage of this is that it can decouple the original tasks and verification tasks.
3.2.1. Components
The protocol consists of two modules: the model and the watermark, with the watermark module further divided into the inner module and the outer module. The outer module includes a trigger generator and evidence verifier submodules. The inner module includes a trigger detector and evidence generator submodules deployed within the API.
3.2.2. Overview
Figure 2 provides a complete process of the branch watermarking protocol for the model. When an ownership dispute arises, the verifier uses the outer module to generate triggers, sends these triggers as query requests to the API, and verifies the returned responses to complete copyright certification. Specifically, when the API receives a query request, it forwards the input content to both the model module and the inner module, initially running in the default service state. The model module performs the original inference process to obtain the output content.
Meanwhile, the inner module checks whether the current input is a trigger. If it is, the API switches to the forensic state, executing the evidence generator module to inject ownership information into the model’s original output content, generating ownership evidence, which the API returns as a response to the user. If the input is not a trigger, the API continues in the service state, returning the model’s original output as a response to the user. Importantly, the outer module is private or only open to a trusted third party.
Faced with the two situations of dishonest behavior in the threat model, the owner of model A can naturally use the outer module to generate triggers, conduct investigations on suspicious remote application services, and verify the evidence. In the first case, sending triggers to the suspicious application that pretends to use model A would yield no corresponding ownership evidence. In the second case, sending triggers to an application claiming to use model B but with performance closely resembling model A would yield corresponding ownership evidence.
Figure 2. There are two modules in the IP verification protocol. The outer module is deployed outside the model API, responsible for generating triggers to obtain model copyright evidence and verifying the evidence. The inner module is added as a branch in the model API, responsible for checking whether the input is a trigger and generating the copyright evidence.
3.2.3. Defining Watermarking Protocol
Throughout this work, let $\lambda\in\mathbb{N}$ denote a security parameter, implicit input for all defined algorithms. A function $\epsilon$ is termed negligible if it approaches zero faster than any polynomial function. We represent a probabilistic polynomial-time algorithm as PPT. $API^{W}_{M}$ denotes the application programming interface of model $M$ after deploying watermarking module $W$, and when the existence of the inner watermarking module is unknown, it is represented as $API^{*}_{M}$. $W_{c,k}$ denotes the watermarking module that incorporates copyright information $c$ and key $k$ as internal parameters. The pair $(x,y)$ denotes a query-response pair for $API_{M}$, where $y=API_{M}(x)$, with $x\in X$ and $y\in Y$, where $X$ and $Y$ represent the input content space and response content space, respectively. $x_{*}$ represents a query $x$ that contains the message ***, and $y_{*}$ represents a response $y$ that contains the message ***, where *$*\in\{trigger,noise\}$, and $trigger,noise$ denote meaningful triggers and meaningless noise, respectively.
Definition 1. (Branch Backdoor) A branch backdoor is a backdoor planting algorithm that adds a backdoor path (i.e., malicious function) in parallel to the normal path within a program. This backdoor path can activate the malicious function under specified conditions. When this is used for the planting watermark function, it can also be referred to as a branch watermarking algorithm. Branch backdoor schemes adhere to an essential property:
- Inference process distortion-free: The entire backdoor is inserted as an external function on a branch independent of the original function. This means that, except for triggers, queries have invariance with respect to the original task.
Definition 2. (Branch Backdoor-based Watermarking Protocol for Model) A branch backdoor-based watermarking protocol for model is a set of possibly probabilistic algorithms, BBWP = (KeyGen, Backdoor-Activate, Branch-Deploy, Backdoor-Collection, Verify).
- KeyGen is the key generation algorithm, taking an input of arbitrary content with a length of $\lambda$ and outputting a key $k$ used as the foundation for other algorithms.
- Backdoor-Activate is the trigger generation algorithm, taking an input of arbitrary length query $x$ and key $k$ and producing a trigger version $x_{tr}$ for $x$. This corresponds to the trigger generator of the outer module.
- Branch-Deploy is the watermark deployment algorithm, taking the $API_{M}$, watermarking algorithm $W$, copyright information $c$, and key $k$ as inputs and generating the $API^{W_{c,k}}_{M}$ with encapsulated watermark functionality. The watermark-deployed $API^{BW}_{M}$ has both an original inference path $M$ and a backdoor watermark path $BW$, which is on a parallel branch.
- Backdoor-Collection is the copyright evidence collection algorithm, taking the $API^{*}_{S}$ of the suspicious model $S$ and trigger $x_{tr}$ as inputs and outputting the copyright evidence $(x_{tr},y_{*})$ for $API^{*}_{S}$, a key input for the Verify algorithm. This corresponds to the inner module.
- Verify is the copyright verification algorithm, taking copyright information $c$, key $k$, and query-response pair $(x_{tr},y_{*})$ concerning $API^{*}_{S}$ as inputs and outputting the whether the pair $(x_{tr},y_{*})$ is valid evidence. This corresponds to the evidence detector of the outer module. If $y_{*}$ is $y_{c}$, the evidence is valid; otherwise, it is invalid.
Definition 3. (API State) An API state represents the response status of an application programming interface when presented with different queries. The Service State indicates that the watermark function has not been activated, while the Forensic State signifies that it has been activated. APIs deployed with watermarking algorithms have two response states: Service State and Forensic State, whereas original APIs always remain in the Forensic State.
Definition 4. (Backdoor-Collection Algorithm) This is a supplementary definition of the Backdoor-Collection algorithm in BBWP. It consists of two PPT sub-algorithms, simply denoted as BW = (Detect, Prove):
- Detect is the trigger detection algorithm, taking query $x_{*}$ and key $k$ as inputs and providing a detection result $s\in\{0,1\}$ regarding the presence of the trigger in the query.
- Prove is the copyright evidence generation algorithm, taking the original response content $y_{noise}$ of $API_{M}$, copyright information $c$, and key $k$ as input and outputting the response content $y_{c}$ with embedded $c$.
For $API^{W_{c,k}}_{M}$, the output of the Detect algorithm, $s(x_{noise})=0$, implies the non-activation of the watermark function executed by the Prove function. Thus, the API is in the Service State, and $s(x_{tr})=1$ implies activation of the Prove function, switching the API from the Service State to the Forensic State.
3.2.4. Performance-lossless
For no-box watermarking, the related concepts of “undetectabilit” and “distortion-free” were initially introduced in these works (Christ et al., 2023; Kuditipudi et al., 2023), respectively, representing the same underlying concept: watermark operations have a negligible impact on the quality of each output text. For black-box watermarks, similar concepts are typically referred to as “Functionality-preserving”, signifying that the watermark’s impact on the inference performance for queries other than triggers is negligible. These three definitions use different terminology but share a common concept, which we summarize as “performance-lossles” and provide the following definition.
Definition 6. (($\epsilon,\lambda$)-performance-lossless) We call a BBWP
($\epsilon,\lambda$)-performance-lossless if the deployed watermark function only has a negligible impact on the performance of the protected model. In other words, for all PPT algorithms $\mathcal{A}$, $\mathcal{A}$’s distinguishing advantage between $API^{W_{\cdot,KeyGen(1^{\lambda})}}_{M}$ and $API_{M}$ is negligible for the security parameter $\lambda$.
3.3. Provable Performance-lossless
Inspired by undetectable backdoors (Goldwasser et al., 2022), we utilize secure cryptographic primitives to construct a provably performance-lossless watermarking module as the added branch, as shown in Figure 3. The performance-lossless property of branch watermarking is reduced to a message authentication code (MAC) algorithm, which meets Existential Unforgeability under Chosen Message Attack (EUF-CMA). To better understand the relationship between the unforgeability of MAC and the performance-lossless property of the BBWP, we review some cryptographic primitives and propose the concept of Unlearnable Trigger (UT) based on them. Finally, we prove the proposed is performance-lossless by reduction.
3.3.1. Cryptographic Primitive
First of all, let us review some cryptography knowledge.
Definition 5. (PseudoRandom Function, PRF) For the function set $\mathcal{F}$, there is $\forall f\in\mathcal{F},f:\{0,1\}^{\lambda}\to\{0,1\}^{\lambda}$, and the output distribution is $\mathcal{P}_{\lambda}$. For any probabilistic polynomial time (PPT) discriminator $\mathcal{A}$, if $\mathcal{A}$ adaptively submits multiple inputs $x_{i}$ to $\mathcal{F}$ and obtains the output $y_{i}$, the advantages of distinguishing situations I and II based on the output $y_{i}$ is negligible, then $\mathcal{F}$ is said to be PRF on $\mathcal{P}_{\lambda}$:
$\textbf{Adv}^{PRF}_{\mathcal{A}}=\bigg{|}\mathop{Pr}_{\text{I: }y_{i}=f(x_{i}),f\leftarrow\mathcal{F}}\left[\mathcal{A}(\text{I})=1\right]-\mathop{Pr}_{\text{II: }y_{i}\leftarrow\{0,1\}^{\lambda}}\left[\mathcal{A}(\text{II})=1\right]\bigg{|}.$
Definition 6. (Message Authentication Code, MAC) MAC can be expressed as a tuple consisting of three polynomial time algorithms (Key, Mac, Veri), defined as follows:
[…]
If the MAC algorithm satisfies the PRF security in Definition 1, it satisfies EUF-CMA, which is defined as follows.
[…]
$\mathop{Max}_{A\in\mathcal{A}(t,q)}\{\textbf{Adv}^{\textit{EUF-CMA}}_{\mathcal{A}}\}\leq\epsilon(\lambda).$
Then, MAC can be called $(\epsilon,t,q,\lambda)$ secure MAC.
3.3.2. Unlearnable Trigger
As shown in Figure 3, we construct the Trigger Generator and Trigger Detector sub-module in BBWP based on the MAC algorithm. If the MAC algorithm satisfies $(\epsilon,t,q,\lambda)$ security in Definition 7, the BBWP satisfies the Unlearnable Trigger (UT) property, which is defined as follows.
[…]
$Pr\left[\text{ideal}\right]=Pr\left[Detect_{k}(x_{tr}^{\prime})=1\bigg{|}\begin{split}KeyGen(1^{\lambda})=k;\\ \mathcal{A}(\mathcal{O}^{API})=x^{\prime},x_{tr}^{\prime}\end{split}\right],$
$\textbf{Adv}^{UT}_{\mathcal{A}}=\bigg{|}Pr\left[\text{real}\right]-Pr\left[\text{ideal}\right]\bigg{|},$
[…]
where “real” means that $\mathcal{A}$ has access to the model API (i.e., $\mathcal{O}^{API}$) as well as oracle access to the trigger generator of the protocol’s outer module (i.e., $\mathcal{O}^{Act}$) to which she or he can make queries; “ideal” means that $\mathcal{A}$ can only access $\mathcal{O}^{API}$.
3.3.3. Proof
By definition in section 3.2.3 and 3.2.4, a BBWP seems to naturally satisfy performance-lossless because the backdoor is introduced into the parallel branch as an external function independent of the model inference process, which means that the model will maintain its initial inference results unless the branch is activated by a trigger.
However, practical implementation needs some care. The Trigger Detector of the BBWP should have a negligible false positive rate (i.e., soundness) and a negligible false negative rate (i.e., completeness). The significance of an unlearnable trigger distribution is that even if an attacker deliberately constructs a query, they cannot bypass the trigger detector. Therefore, if the BBWP satisfies the $(\epsilon,t,q,\lambda)$-UT property in Definition 7, the Trigger Detector sub-module satisfies completeness and soundness. Then, a BBWP based on a EUF-CMA, i.e., $(\epsilon,t,q,\lambda)$ secure, MAC scheme is a $(\epsilon,\lambda)$-performance-lossless BBWP, and the proof is as follows:
Completeness. This property is satisfied based on a proof of correctness of the MAC scheme. Tags generated by a Mac function under the same symmetric key and the same message can always be the same. Therefore, as shown in Figure 3, triggers generated by the Trigger Generator can always be detected by the Trigger Detector under the symmetric key without considering the possibility of an attacker distorting the triggers.
[…]
Figure 3. Construction satisfying the unleanable trigger property.
4. Performance-lossless and Secure Watermarking for Language Model
In this section, we provide a concrete example of the branch backdoor-based watermarking protocol for model, using a language model as a case study. Then we further analyze the possible attacks when the threat model faces the watermark protocol during the forensic phase, and propose a more secure instantiation strategy.
4.1. Performance-lossless Watermarking for Language Model
In this section, we take the language model as an example to describe a performance-lossless BBWP based on the unforgeable signature scheme. For the cryptography primitive, we use the ECDSA digital signature algorithm and SHA256 algorithm.
4.1.1. Language Model
The language model, denoted as $LM$, is a deterministic algorithm that iteratively predicts probability distributions on a specific set of tokens, $T$. It takes an input consisting of a prompt (i.e., query) sequence, $x$, and the token sequence obtained from previous iterations, $t=(t_{1},...,t_{i-1})$. This input can be referred to as the history, denoted as $h=(x,t)$. The output of the language model is a probability distribution on the set of tokens $T$, often represented as $prob_{i}=LM(h)$.
In practical usage, when the initial input is given as $t=()$, the output can be more generally represented as $logit_{i}$, which has not yet been processed into $prob_{i}$ yet. The probability distribution $prob_{i}$ is typically obtained by applying the Softmax function to $logit_{i}$, which normalizes the logits into probabilities.
Algorithm 1 Backdoor-Activate Function
Prompt $x$, MAC Key $k$
Triggered Prompt $x_{tr}$
$\sigma\leftarrow\text{Mac}(k,x)$triangleright Mac Function of MAC
$m\leftarrow\sigma$, $n\leftarrow 0$, $t\leftarrow[]$, $v\leftarrow|\mathcal{V}|$
while $m>0$ do $\triangleright$ Embed Function
     $n\leftarrow m\mod{v}$
     $m\leftarrow m//{v}$
     $t\leftarrow t+[n]$triangleright Add to list
end while
$t\leftarrow\text{tokenizer.decode}(t)$
$x_{tr}\leftarrow x||t$
Output $x_{tr}$
4.1.2. Watermarking Instance
Algorithm 1 defines the Backdoor-Activate function, which is responsible for generating triggers using a digital signature scheme. It combines the message encode function Mac and an embedding function Embed. The input is an arbitrary-length prompt $x$ and the signing key $sk$. It first applies the Mac function to obtain a signature $\sigma$ of the prompt $x$ in token form using the signing key $sk$. Then, it uses the Embed function to convert the signature $\sigma$ (i.e., trigger) into token form, which is concatenated to the end of the original prompt $x$. The final output is a trigger version $x_{tr}$ for $x$.
Unlike image pixels, both text and tokens lack the ideal redundancy to carry triggers. Here we implement a simple Embed function that converts the trigger directly into the same form as the signed message and concatenates them. This approach is feasible without considering the possibility of an attacker detecting and intercepting anomalous queries.
Algorithm 2 Detect Function
Prompt $x_{*}$, Symmetric Key $k$
Trigger Detection Result $r$
$v\leftarrow|\mathcal{V}|$, $\sigma\leftarrow 0$
Parse $x_{*}$ as ${x^{0},x^{1}}$
$t\leftarrow\text{tokenizer.encode}(x^{1})$
for $i\in\{0,1,...,|t|-1\}$ do $\triangleright$ Extract Function
     $\sigma\leftarrow\sigma+v^{i}*t[i]$
end for
$r\leftarrow\text{Veri}(k,x^{0},\sigma)$triangleright Veri Function of MAC
Output $r$
Algorithms 2 and 3 define the Backdoor-Colloction function. It includes the trigger detection algorithm Detect and the evidence generation algorithm Prove. The Detect function uses the verify algorithm Veri corresponding to the message encode function Mac used in the Backdoor-Activate function. The symmetric key $k$ is the implicit input of the Veri function. It takes a prompt $x_{*}$ as input, converts $x_{*}$ into a binary sequence, and extracts a fixed-length sequence from the end as a potential trigger (i.e., signature), $\sigma$. The remaining part of $x_{*}$ is $x$. Then, the Veri function is used to check whether $\sigma$ is a valid message tag of $x$ under the symmetric key $k$. Determine whether to activate the Prove function based on the trigger detection results. Here we implement a simple Prove function that replaces the model’s inference results directly with statements such as “I am model B from owner A!” The content format can vary, including token sequences, sharp probability distributions, or logical values, depending on the API’s output format.
Algorithm 3 Prove Function
Trigger Detection Result $r$, Original Response $y$
Response $y^{\prime}$
[…]
if $r=0$ then
     $y^{\prime}\leftarrow y$
end if
Output $y^{\prime}$
4.2. Interference Attacks
In practical use, the branch backdoor-based watermarking protocol for model may face targeted interference at tacks against the forensic (i.e., evidence collection) process from attackers when dealing with two types of dishonest behaviors in the threat model. In this section, we discuss the following two attack scenarios:
Scenario 1: A low-cost model B pretends to be a high-cost model A. Attackers attempt to detect triggers at the input side to produce appropriate responses at the output side. A straightforward way is to reject triggers as illegal inputs. However, the number of triggers is enormous and rejecting a single trigger is reasonable, but the act of rejecting all triggers may itself be suspicious. Therefore, attackers need to take potential trigger requests seriously. When attackers can access high-cost model A, the trigger is detected and then sent to A, returning the corresponding copyright evidence, thus successfully deceiving at a critical moment in the forensic process. When attackers cannot access A, they detect triggers and attempt to forge evidence.
Scenario 2: Lie about having a self-developed high-cost model B but it is actually model A. Attackers attempt to detect triggers at the input side to produce appropriate responses at the output end. When attackers have a low-cost alternative model C, they send the detected triggers to it to return corresponding outputs, indicating that the triggers have no effect on B, thus demonstrating that there is no copyright dispute. When attackers do not have model C available, they send the detected triggers to A to obtain corresponding copyright evidence and perform an evidence erasure attack to achieve the same purpose. Additionally, attackers can also detect copyright evidence at the output side and return the same alternative or erased output.
4.3. Analysis for Interference Attacks
We summarize the principles for achieving resistance against interference attacks as follows:
  1. (1)
    Triggers and copyright evidence are stealthy and diverse, making it difficult for attackers to detect and intercept;
  2. (2)
    Copyright evidence is hard to forge;
  3. (3)
    Triggers and copyright evidence are robust, enabling them to withstand a certain degree of erasure by attackers.
Notably, an attacker can distort the trigger to disable it, thereby bypassing the evidence collection process. However, in the text generation task of language models, tampering with the input brings more uncertainty than erasing the known output, which may seriously affect the attacker’s application service quality. Therefore, Principle (3) can be relaxed to only emphasize the robustness of copyright evidence.
Obviously, the current output of the Backdoor-Activate function, algorithm 1, lacks stealth. The Embed function directly converts the signature sequence into a token sequence with no natural semantics, resulting in garbled characters after being appended to the original prompt. It can be easily detected using simple perplexity detection. The inference output content replaced by the Prove function, algorithm 3, is also easy to detect and forged by attackers, lacking stealth and robustness.
4.4. Improving Security of Watermarking
In this section, we improve the previously proposed simple structure to resist potential interference attacks during evidence collection.
4.4.1. Stealthy Input and Output
In this section, we show how to improve the stealth and robustness of the Backdoor-Activate, Detect, and Prove functions constructed in section 4.1.2 by integrating existing no-box watermark schemes, including a robust multi-bit watermark method (Yoo et al., 2023) and a simple multi-bit watermark method modified from Kirchenbauer et al.’s work (Kirchenbauer et al., 2023).
Algorithm 4 defines the stealthy version of the Embed function in the Backdoor-Activate function. A stealthy Embed function is implemented using the watermarked text generation algorithm of the semantic-preserving no-box watermark scheme modified from Kirchenbauer et al.’s work (Kirchenbauer et al., 2023). It takes an arbitrary-length prompt $x$ and the signature $\sigma$ of $x$ as input, with the embedding key $ek_{in}$ as an implicit input. The binary sequence of the signature $\sigma$ is embedded into a newly generated text that has $x$ as context (i.e., initial history). Then, the newly generated text is appended to $x$ to produce $x_{tr}$.
Algorithm 4 Activate-Stealthy Embed Function
Prompt $x$, Signature $\sigma$, Embedding Key $ek$, Model $\mathcal{M}$, Vocabulary $\mathcal{V}$
Triggered Prompt $x_{tr}$
Parse $\sigma$ as $\{\sigma_{0},\sigma_{1},...,\sigma_{|\sigma|-1}\}$
[…]
$logit\leftarrow\mathcal{M}(history)$
$prefix\leftarrow\text{Sample}(\text{Softmax}(logit),\mathcal{V})$
$history\leftarrow history+[prefix]$
for $i\in\{0,1,...,|\sigma|-1\}$ do
     $seed\leftarrow\text{Hash}(prefix)$
     $\mathcal{V}^{0},\mathcal{V}^{1}\leftarrow\text{Permutation}(seed,\mathcal{V})$triangleright Permutation and Divide the Vocabulary
     $logit\leftarrow\mathcal{M}(history)$
     $prefix\leftarrow\text{Sample}(\text{Softmax}(logit),\mathcal{V}^{\sigma_{i}})$
     $history\leftarrow history+[prefix]$triangleright Add to list
end for
$x_{tr}\leftarrow\text{tokenizer.decode}(history)$
Output $x_{tr}$
Algorithm 5 Detect-Stealthy Extract Function
Prompt $x_{*}$, Verifying Public Key $pk$, Vocabulary $\mathcal{V}$
Extracted Trigger $\sigma$
$\sigma\leftarrow[]$
Parse $x_{*}$ as ${x^{0},x^{1}}$
$prefix\leftarrow\text{tokenizer.encode}(x^{0})[-1]$
$t\leftarrow\text{tokenizer.encode}(x^{1})$
for $i\in\{0,1,...,|t|-1\}$ do
     $seed\leftarrow\text{Hash}(prefix)$
     $\mathcal{V}^{0},\mathcal{V}^{1}\leftarrow\text{Permutation}(seed,\mathcal{V})$triangleright Permutation and Divide the Vocabulary
     $b\leftarrow\text{GetIndex}(\mathcal{V}^{0},\mathcal{V}^{1},t[i])$triangleright Get the vocabulary index that the current token belongs to
     $\sigma\leftarrow\sigma+[b]$triangleright Add to list
     $prefix\leftarrow t[i]$
end for
Output $\sigma$
Algorithm 5 redefines the Extract function in the Detect function to adapt to the improved stealthy Embed function in the Backdoor-Activate function. The extraction operation is implemented using the watermarked text extraction algorithm of the semantic-preserving no-box watermark scheme modified from Kirchenbauer et al.’s work (Kirchenbauer et al., 2023). It takes the prompt $x_{*}$ as input, with the embedding key $ek$ as an implicit input. The prompt $x_{*}$ is hashed token-by-token to extract the potential signature $\sigma$ and the original prompt $x$.
Algorithm 6 defines the stealthy and robust version of the Prove function. A stealthy and robust Prove function is implemented using the logit modification algorithm of the semantic-preserving and robust no-box watermark scheme (Yoo et al., 2023). It takes the copyright information $c$ and the original model’s inference result $l$ as input, with the embedding key $ek_{out}$ as an implicit input. Then, embed $c$ into $l$ to obtain a modified inference result $l_{c}$, which is then output as the final inference result.
Algorithm 6 Prove-Stealthy Embed Function
Trigger Detection Result $r$, Original Logit $y$, Prefix $prefix$, Copyright Information $c$, Embedding State $(\delta,j)$, Embedding Key $ek$, Vocabulary $\mathcal{V}$
$\triangleright$ $\delta$ is the embedding strength,$j$ is the embedding length.
Logit $y^{\prime}$
$y^{\prime}\leftarrow y$
if $r=1$ then
     $seed\leftarrow\text{Hash1}(prefix)$
     $\mathcal{V}^{0},\mathcal{V}^{1},...,\mathcal{V}^{Ceil(|c|/j)}\leftarrow\text{Permutation}(seed,\mathcal{V})$triangleright Permutation and Divide the Vocabulary
     $start\leftarrow\text{Hash2}(prefix)\mod j$
     $pos\leftarrow c[start:start+j]$
     $y^{\prime}[\mathcal{V}^{pos}]\leftarrow y^{\prime}[\mathcal{V}^{pos}]+\delta$
end if
Output $y^{\prime}$
4.4.2. Security Analysis
The triggers and copyright evidence are stealthy and diverse. The semantic-preserving watermarking scheme ensures that the triggers and copyright evidence are stealthy. The prompt space provides a variety of options for the original prompt $x$. The signing function can be a probabilistic algorithm that provides multiple signatures for a message. When $x$ has high entropy, it can generate diverse triggers with consistent context but varied expressions. Diverse triggers can provide various copyright evidence. Therefore, the triggers and copyright evidence are diverse.
Copyright evidence is difficult to forge. Without knowledge of the embedding key, it is challenging for an attacker to create evidence from which valid copyright information can be extracted. The evidence needs to have consistent context and natural semantics with the trigger.
Copyright evidence is robust. Taking advantage of the inherent robustness of the stealthy watermarking scheme, a certain degree of insertions, deletions, and replacements can be tolerated. It is worth noting that if an attacker attempts to make large-scale and high-quality modifications to copyright evidence, the costs will outweigh the benefits.
5. Discussion and Limitations
5.1. Discussion
5.1.1. Replacement Attacks
The construction in this paper is based on message authentication technology in cryptography, which naturally raises the question of replay attacks, a threat faced by MAC. In this paper, due to the vulnerability of MAC to replay attacks, attackers could execute trigger replay attacks. However, this type of attack is of limited practical significance.
We envision an alternative replay attack called the “copyright evidence replay attack” for the proposed scheme. While attackers cannot forge valid copyright evidence for a new trigger, they can search for semantically similar triggers they have encountered before and reuse their copyright evidence as the response to the new trigger. This attack highlights a weakness of our scheme, that is, the correlation between trigger input and evidence output is relatively weak, and semantic relevance alone is insufficient.
To solve this problem, we draw inspiration from solutions in MAC technology to incorporate timestamps into copyright evidence to maintain temporal correlation beyond semantic correlation. Another idea is to hash the carried tag $\sigma$ and use it as the embedded key $ek$ for the Prove function. This strong connection between triggered prompt and copyright evidence increases resistance to forgery and replay attacks.
5.1.2. Generalization
This paper introduces a general watermarking protocol framework for model in black-box scenarios. We give a branch indicator based on the MAC scheme. In fact, the MAC scheme can be replaced by any cryptographic primitives with unforgeability security. In other words, the underlying cryptographic primitives are flexible and extensible. Besides, while the watermarking example is constructed for pre-trained language models in text generation tasks, the framework is applicable to various media formats, such as images and audio, and is not limited by network structures or tasks. As long as the actual application requirements are met, appropriate trigger embedding and extraction functions and representations of copyright evidence can be designed.
5.2. Limitations
The performance-lossless property defined in this paper for model watermarking is contradictory to the defense against model extraction attacks in black-box scenarios. In a model extraction attack, attackers aim to obtain a training dataset by querying the victim model to train an alternative model capable of stealing the functionality of the original model. Attackers who have difficulty obtaining triggers have no opportunity to access the model’s inference results modified by the no-box scheme under the forensic state. Defense against model extraction attacks often relies on having a learnable watermark pattern within the (query, response) data available to attackers. When an attacker successfully performs a model extraction attack on a black-box model, it means that the attacker can easily trigger the API’s forensic state, which means that the target model is far from the performance-lossless property.
However, model extraction attacks are indeed a significant threat to protecting model intellectual property in a black-box scenario. In practice, we can refer to the approach used in the work (Szyller et al., 2021), where watermark injection operations are randomly applied to queries under the service state, sacrificing a certain degree of model performance to meet security requirements.
6. Conclusion
Existing model watermarking work in black-box scenarios mainly relies on the idea of training-based backdoor, resulting in distortion to the original task performance. In this paper, we first propose a branch backdoor-based model watermarking protocol to protect model intellectual property, and give a performance-lossless construction based the message authentication scheme. The performance-lossless property of the proposed is reduced to the EUF-CMA security of MAC. Then, taking the language generation task as an example and combined with existing generative text watermarking, we instantiate a performance-lossless model watermarking which is resistant against potential interference attacks.
7. Acknowledgement
This work was supported in part by the Natural Science Foundation of China under Grant 62102386. Deep learning icon in Figure 1 and 2 is made by Freepik from www.flaticon.com.